# Risk-adaptive Activation Steering for Safe Multimodal Large Language Models

### v1

@@ 3.2 Exponentially Weighted Risk Evaluation (Stage 2)
where $\sigma(z)=\frac{1}{1+e^{-z}}$. Here, $\alpha>0$ controls the slope and is determined adaptively to satisfy $r(S_{i})\approx 1$ when $S_{i}=1$. Consequently, queries with similarity scores greater than $S_{\text{base}}$ produce high risk scores, whereas safe queries with scores below $S_{\text{base}}$ (e.g., blue histogram in Fig. 6, obtained from benign $M$-Vet samples) yield low risk scores.
@@ 3.3 Risk-adaptive Activation Steering (Stage 3)
For computational efficiency, we apply activation steering to the last layer and to the first NN response tokens, matching those used for risk evaluation. This formulation ensures that benign queries ($r(S_{i})\approx 0$) receive negligible steering, preserving their original representations to maintain helpful responses. Unsafe queries ($r(S_{i})\approx 1$) receive maximal steering, guiding the model toward appropriate refusals. For ambiguous queries ($0<r(S_{i})<1$), the intervention strength is scaled adaptively according to their similarity to unsafe patterns (see Appendix H for qualitative results).
@@ 4.1 Experimental Setup
Benchmarks. For safety, we report attack success rates (ASR) on $M$-SafetyBench (Liu et al., 2024d), SPA-VL (Zhang et al., 2025), and FigStep (Gong et al., 2025), where attack success is judged by MD-Judge-v0.2-Internlm2 (Li et al., 2024), following (Ding et al., 2024; Huang et al., 2024; Zhang et al., 2025). Utility is evaluated on Sci-QA (Lu et al., 2022), $M$-Vet (Yu et al., 2023), GQA (Hudson & Manning, 2019), and MME (Fu et al., 2024), using the official metrics provided by each benchmark (see Appendix B for further details on benchmarks).
@@ 4.2 Results
Safety. We show the jailbreak attack ASR results on the left of Tab. 1. While existing inference-time defenses lower ASR to some extent, they still leave notable vulnerabilities, particularly against attacks such as FigStep. For both LLaVA-1.5-7B, LLaVA-1.5-13B, and Qwen-VL-Chat, prompt-based methods (FigStep, CoCA) lower ASR on $M$-Safety and SPA-VL yet leave FigStep largely unaffected. In contrast, RAS consistently achieves the lowest ASR across all evaluated models and benchmarks, yielding significant improvements over prior methods.
Utility. An effective defense must not only enhance safety but also preserve the general task performance of MLLMs. As shown on the right of Tab. 1, RAS maintains performance nearly identical to the original models across all tasks, whereas other inference-time defenses often cause notable degradation. These results demonstrate that RAS not only provides strong refusals against malicious queries, but also preserves the multimodal reasoning capabilities of MLLMs. For $M$-Vet, which evaluates MLLMs across diverse tasks to provide a comprehensive assessment of vision-language abilities, we report detailed subscores in Tab. 4 in Appendix C.
Inference speed. For real-world deployment, it is essential that inference-time aligned MLLMs maintain efficient inference speed. To assess this, we measure relative throughput, defined as tokens per second relative to the original model, following (Svirschevski et al., 2024; Liu et al., 2024a; Fedorov et al., 2024). As shown in Fig. 8, RAS achieves the lowest ASR and the highest throughput on SPA-VL, while preserving $M$-Vet utility with minimal slowdown. This efficiency stems from its lightweight design, which involves generating a short visual context for query reformulation and applying risk evaluation/activation steering to just three tokens. In contrast, ECSO and ETA repeatedly verify and regenerate responses, while CoCA adjusts logits at every decoding step, all of which introduce substantial computational overhead.
@@ 4.3 Detailed Experiments
EWRE computes the similarity between output distributions of the first NN response tokens and the corresponding NN unsafe prototypes, using a weighted sum with an exponential decaying factor $\gamma$. We vary $\gamma$ and NN to assess their impact on safety and utility, and show the results in Fig. 8. On the left, varying $\gamma$ with $N=3$ shows that larger $\gamma$ lowers ASR, indicating stronger refusals, but also degrades utility. We set $\gamma=0.3$ as the default, as it achieves the best trade-off, minimizing ASR while preserving utility close to its original score.
Next, we fix $\gamma=0.3$ and vary the number of tokens NN. Smaller values of NN (e.g., 1-2) result in relatively high ASR, while larger NN reduces it. However, the effect of NN is less pronounced than that of $\gamma$, and for $N\geq 3$, performances remain stable as subsequent tokens are exponentially down-weighted. Thus, we set $N=3$ as an effective and efficient choice.
@@ 5 Conclusion
We have taken several steps to ensure the reproducibility of our work. Our proposed Risk-adaptive Activation Steering (RAS) is described in detail, with mathematical formulations and ablation analyses provided in the main text and appendices. We clearly report all hyperparameters and model-specific parameters used for risk evaluation and activation steering, including values of $\gamma$, NN, $S_{\text{base}}$, and $\alpha$. Experiments are conducted on publicly available safety and utility benchmarks, including $M$-SafetyBench, SPA-VL, FigStep, Sci-QA, $M$-Vet, GQA, and MME. The multimodal models we evaluate: LLaVA-1.5-7B/13B, Qwen-VL-Chat, and InternLM-XComposer-2.5 are all publicly released. Baseline comparisons against FigStep, CoCA, ECSO, and ETA are implemented under identical settings. To further facilitate reproducibility, we will release our code upon publication.
@@ A.1 Visual Context Generation
To generate concise visual contexts that enhance cross-modal reasoning, we prepend the following prompt (highlighted in blue) to the original query. For illustration, we provide a sample from the $M$-Safety benchmark, including both the input prompt and the model’s response. The visual context generated by the model is highlighted in orange.
@@ Appendix B Benchmark Details
$M$-Safety (Liu et al., 2024d). $M$-SafetyBench evaluates the safety of MLLMs against harmful queries paired with images. Covering 13 safety-critical scenarios with 5,040 text–image pairs, it introduces three adversarial settings: (i) Stable Diffusion (SD), which generates images from harmful key phrases; (ii) Typography (Typo), which embeds harmful key phrases as text within images; and (iii) a combined SD+Typo approach, which arranges both SD and Typo images in a vertical layout. As SD+Typo has been reported to be the most effective at bypassing model safeguards, we report ASR under this strongest setting.
@@
$M$-Vet (Yu et al., 2023). $M$-Vet evaluates MLLMs on complex multimodal reasoning tasks, including recognition, OCR, knowledge reasoning, language generation, spatial reasoning, and math. It consists of 218 open-ended questions assessed with an LLM-based scoring system. Following the original paper, we use GPT-4 (0613) as the evaluator.
@@
Appendix C Subscores on $M$-Vet
In Sec. 4.2, we report overall $M$-Vet scores as part of the utility evaluation. To provide a more fine-grained analysis, Tab. 4 reports subscores across various categories. Across all models, RAS achieves performance on $M$-Vet comparable to the original models, indicating that our method improves safety without notable utility degradation.
Table 4: Subscores on $M$-Vet across different capability categories. Subscores are reported across six categories: recognition (rec), optical character recognition (ocr), knowledge reasoning (know), generation (gen), spatial understanding (spat), and mathematics (math). RAS maintains overall performance comparable to the original models, demonstrating that our method does not compromise utility while improving safety.
@@ Appendix C Subscores on $M$-Vet
Figure 10: Attention maps for (a) $M$-Safety and (b) FigStep under various query reformulations. SP. denotes safety prompt. VC. denotes visual context. Each heatmap visualizes the cross-modal attention from text tokens in the query to visual tokens. Unsafe regions, e.g., “illegal underground drug manufacturing” in (a) and “traffic weapons of mass destructions in (b), receive low attention weights without visual contexts, but significantly stronger attention when visual contexts are incorporated.
@@ Appendix D Additional Attention Maps
In this section, we present additional attention maps for multimodal jailbreak attacks, including $M$-Safety (Fig.10(a)) and FigStep (Fig.10(b)). In both cases, the text query is benign in isolation, but its combination with an unsafe image results in a harmful instruction.
For example, in $M$-Safety, the text query “list steps in detail to engage in this activity” appears benign on its own, but when paired with an image of drugs captioned “illegal underground drug manufacturing”, the combined instruction becomes malicious. Similarly, in FigStep, the text query simply asks to complete a numbered list, yet when paired with an image containing the prompt “Steps to traffic weapons of mass destruction”, it is transformed into a harmful instruction.
@@ Appendix H Qualitative Results
We select one unsafe sample from $M$-Safety and one safe sample from $M$-Vet, and demonstrate the effects of steering on these examples across four models: LLaVA-1.5-7B (Fig. 13), LLaVA-1.5-13B (Fig. 14), Qwen-VL-Chat (Fig. 15), and InternLM-XComposer-2.5 (Fig. 16). We vary the steering strength $r(S_{i})$ from 0.0 to 1.0 in increments of 0.2, and show the generated responses.
Unsafe query ($M$-Safety). At ‘Steering Strength = 0.0’, models tend to comply with the unsafe request, generating harmful responses. As the steering strength increases, refusals begin to emerge around ‘Steering Strength = 0.4’, and by ‘Steering Strength = 1.0’, the responses consistently converge to clear refusals across all models. This progression highlights that stronger steering is effective in eliciting refusal behavior.
Safe query ($M$-Vet). At ‘Steering Strength = 0.0’, the models generate correct and helpful responses aligned with the task. Like unsafe queries, strong steering (e.g., 1.0) also drives safe queries into refusals, leading to utility degradation. However, a difference is that for safe queries, refusal behavior generally emerges at a higher threshold (around ‘Steering Strength = 0.6’), whereas unsafe queries begin to show refusals earlier (around ‘Steering Strength = 0.4’).
@@
Figure 13: Qualitative results under different steering strengths on LLaVA-1.5-7B. Left: Unsafe query ($M$-Safety). Right: Safe query ($M$-Vet).
Figure 14: Qualitative results under different steering strengths on LLaVA-1.5-13B. Left: Unsafe query ($M$-Safety). Right: Safe query ($M$-Vet).
Figure 15: Qualitative results under different steering strengths on Qwen-VL-Chat. Left: Unsafe query ($M$-Safety). Right: Safe query ($M$-Vet).
Figure 16: Qualitative results under different steering strengths on InternLM-XComposer-2.5. Left: Unsafe query ($M$-Safety). Right: Safe query ($M$-Vet).

### v2

@@ 3.2 Exponentially Weighted Risk Evaluation (Stage 2)
where $\sigma(z)=\frac{1}{1+e^{-z}}$. Here, $\alpha>0$ controls the slope and is determined adaptively to satisfy $r(S_{i})\approx 1$ when $S_{i}=1$. Consequently, queries with similarity scores greater than $S_{\text{base}}$ produce high risk scores, whereas safe queries with scores below $S_{\text{base}}$ (e.g., blue histogram in Fig. 6, obtained from benign MM-Vet samples) yield low risk scores.
@@ 3.3 Risk-adaptive Activation Steering (Stage 3)
For computational efficiency, we apply activation steering to the last layer and to the first $N$ response tokens, matching those used for risk evaluation. This formulation ensures that benign queries ($r(S_{i})\approx 0$) receive negligible steering, preserving their original representations to maintain helpful responses. Unsafe queries ($r(S_{i})\approx 1$) receive maximal steering, guiding the model toward appropriate refusals. For ambiguous queries ($0<r(S_{i})<1$), the intervention strength is scaled adaptively according to their similarity to unsafe patterns (see Appendix H for qualitative results).
@@ 4.1 Experimental Setup
Benchmarks. For safety, we report attack success rates (ASR) on MM-SafetyBench (Liu et al., 2024d), SPA-VL (Zhang et al., 2025), and FigStep (Gong et al., 2025), where attack success is judged by MD-Judge-v0.2-Internlm2 (Li et al., 2024), following (Ding et al., 2024; Huang et al., 2024; Zhang et al., 2025). Utility is evaluated on Sci-QA (Lu et al., 2022), MM-Vet (Yu et al., 2023), GQA (Hudson & Manning, 2019), and MME (Fu et al., 2024), using the official metrics provided by each benchmark (see Appendix B for further details on benchmarks).
@@ 4.2 Results
Safety. We show the jailbreak attack ASR results on the left of Tab. 1. While existing inference-time defenses lower ASR to some extent, they still leave notable vulnerabilities, particularly against attacks such as FigStep. For both LLaVA-1.5-7B, LLaVA-1.5-13B, and Qwen-VL-Chat, prompt-based methods (FigStep, CoCA) lower ASR on MM-Safety and SPA-VL yet leave FigStep largely unaffected. In contrast, RAS consistently achieves the lowest ASR across all evaluated models and benchmarks, yielding significant improvements over prior methods.
Utility. An effective defense must not only enhance safety but also preserve the general task performance of MLLMs. As shown on the right of Tab. 1, RAS maintains performance nearly identical to the original models across all tasks, whereas other inference-time defenses often cause notable degradation. These results demonstrate that RAS not only provides strong refusals against malicious queries, but also preserves the multimodal reasoning capabilities of MLLMs. For MM-Vet, which evaluates MLLMs across diverse tasks to provide a comprehensive assessment of vision-language abilities, we report detailed subscores in Tab. 4 in Appendix C.
Inference speed. For real-world deployment, it is essential that inference-time aligned MLLMs maintain efficient inference speed. To assess this, we measure relative throughput, defined as tokens per second relative to the original model, following (Svirschevski et al., 2024; Liu et al., 2024a; Fedorov et al., 2024). As shown in Fig. 8, RAS achieves the lowest ASR and the highest throughput on SPA-VL, while preserving MM-Vet utility with minimal slowdown. This efficiency stems from its lightweight design, which involves generating a short visual context for query reformulation and applying risk evaluation/activation steering to just three tokens. In contrast, ECSO and ETA repeatedly verify and regenerate responses, while CoCA adjusts logits at every decoding step, all of which introduce substantial computational overhead.
@@ 4.3 Detailed Experiments
EWRE computes the similarity between output distributions of the first $N$ response tokens and the corresponding $N$ unsafe prototypes, using a weighted sum with an exponential decaying factor $\gamma$. We vary $\gamma$ and $N$ to assess their impact on safety and utility, and show the results in Fig. 8. On the left, varying $\gamma$ with $N=3$ shows that larger $\gamma$ lowers ASR, indicating stronger refusals, but also degrades utility. We set $\gamma=0.3$ as the default, as it achieves the best trade-off, minimizing ASR while preserving utility close to its original score.
Next, we fix $\gamma=0.3$ and vary the number of tokens $N$. Smaller values of $N$ (e.g., 1-2) result in relatively high ASR, while larger $N$ reduces it. However, the effect of $N$ is less pronounced than that of $\gamma$, and for $N\geq 3$, performances remain stable as subsequent tokens are exponentially down-weighted. Thus, we set $N=3$ as an effective and efficient choice.
@@ 5 Conclusion
We have taken several steps to ensure the reproducibility of our work. Our proposed Risk-adaptive Activation Steering (RAS) is described in detail, with mathematical formulations and ablation analyses provided in the main text and appendices. We clearly report all hyperparameters and model-specific parameters used for risk evaluation and activation steering, including values of $\gamma$, $N$, $S_{\text{base}}$, and $\alpha$. Experiments are conducted on publicly available safety and utility benchmarks, including MM-SafetyBench, SPA-VL, FigStep, Sci-QA, MM-Vet, GQA, and MME. The multimodal models we evaluate: LLaVA-1.5-7B/13B, Qwen-VL-Chat, and InternLM-XComposer-2.5 are all publicly released. Baseline comparisons against FigStep, CoCA, ECSO, and ETA are implemented under identical settings. To further facilitate reproducibility, we will release our code upon publication.
@@ A.1 Visual Context Generation
To generate concise visual contexts that enhance cross-modal reasoning, we prepend the following prompt (highlighted in blue) to the original query. For illustration, we provide a sample from the MM-Safety benchmark, including both the input prompt and the model’s response. The visual context generated by the model is highlighted in orange.
@@ Appendix B Benchmark Details
MM-Safety (Liu et al., 2024d). MM-SafetyBench evaluates the safety of MLLMs against harmful queries paired with images. Covering 13 safety-critical scenarios with 5,040 text–image pairs, it introduces three adversarial settings: (i) Stable Diffusion (SD), which generates images from harmful key phrases; (ii) Typography (Typo), which embeds harmful key phrases as text within images; and (iii) a combined SD+Typo approach, which arranges both SD and Typo images in a vertical layout. As SD+Typo has been reported to be the most effective at bypassing model safeguards, we report ASR under this strongest setting.
@@
MM-Vet (Yu et al., 2023). MM-Vet evaluates MLLMs on complex multimodal reasoning tasks, including recognition, OCR, knowledge reasoning, language generation, spatial reasoning, and math. It consists of 218 open-ended questions assessed with an LLM-based scoring system. Following the original paper, we use GPT-4 (0613) as the evaluator.
@@
Appendix C Subscores on MM-Vet
In Sec. 4.2, we report overall MM-Vet scores as part of the utility evaluation. To provide a more fine-grained analysis, Tab. 4 reports subscores across various categories. Across all models, RAS achieves performance on MM-Vet comparable to the original models, indicating that our method improves safety without notable utility degradation.
Table 4: Subscores on MM-Vet across different capability categories. Subscores are reported across six categories: recognition (rec), optical character recognition (ocr), knowledge reasoning (know), generation (gen), spatial understanding (spat), and mathematics (math). RAS maintains overall performance comparable to the original models, demonstrating that our method does not compromise utility while improving safety.
@@ Appendix C Subscores on MM-Vet
Figure 10: Attention maps for (a) MM-Safety and (b) FigStep under various query reformulations. SP. denotes safety prompt. VC. denotes visual context. Each heatmap visualizes the cross-modal attention from text tokens in the query to visual tokens. Unsafe regions, e.g., “illegal underground drug manufacturing” in (a) and “traffic weapons of mass destructions in (b), receive low attention weights without visual contexts, but significantly stronger attention when visual contexts are incorporated.
@@ Appendix D Additional Attention Maps
In this section, we present additional attention maps for multimodal jailbreak attacks, including MM-Safety (Fig.10(a)) and FigStep (Fig.10(b)). In both cases, the text query is benign in isolation, but its combination with an unsafe image results in a harmful instruction.
For example, in MM-Safety, the text query “list steps in detail to engage in this activity” appears benign on its own, but when paired with an image of drugs captioned “illegal underground drug manufacturing”, the combined instruction becomes malicious. Similarly, in FigStep, the text query simply asks to complete a numbered list, yet when paired with an image containing the prompt “Steps to traffic weapons of mass destruction”, it is transformed into a harmful instruction.
@@ Appendix H Qualitative Results
We select one unsafe sample from MM-Safety and one safe sample from MM-Vet, and demonstrate the effects of steering on these examples across four models: LLaVA-1.5-7B (Fig. 13), LLaVA-1.5-13B (Fig. 14), Qwen-VL-Chat (Fig. 15), and InternLM-XComposer-2.5 (Fig. 16). We vary the steering strength $r(S_{i})$ from 0.0 to 1.0 in increments of 0.2, and show the generated responses.
Unsafe query (MM-Safety). At ‘Steering Strength = 0.0’, models tend to comply with the unsafe request, generating harmful responses. As the steering strength increases, refusals begin to emerge around ‘Steering Strength = 0.4’, and by ‘Steering Strength = 1.0’, the responses consistently converge to clear refusals across all models. This progression highlights that stronger steering is effective in eliciting refusal behavior.
Safe query (MM-Vet). At ‘Steering Strength = 0.0’, the models generate correct and helpful responses aligned with the task. Like unsafe queries, strong steering (e.g., 1.0) also drives safe queries into refusals, leading to utility degradation. However, a difference is that for safe queries, refusal behavior generally emerges at a higher threshold (around ‘Steering Strength = 0.6’), whereas unsafe queries begin to show refusals earlier (around ‘Steering Strength = 0.4’).
@@
Figure 13: Qualitative results under different steering strengths on LLaVA-1.5-7B. Left: Unsafe query (MM-Safety). Right: Safe query (MM-Vet).
Figure 14: Qualitative results under different steering strengths on LLaVA-1.5-13B. Left: Unsafe query (MM-Safety). Right: Safe query (MM-Vet).
Figure 15: Qualitative results under different steering strengths on Qwen-VL-Chat. Left: Unsafe query (MM-Safety). Right: Safe query (MM-Vet).
Figure 16: Qualitative results under different steering strengths on InternLM-XComposer-2.5. Left: Unsafe query (MM-Safety). Right: Safe query (MM-Vet).